Grade 9Algebra

What is the highest power of 7 that can divide 5000! without leaving a reminder?
Please provide solution.

Subham Chakravorty
8 Years agoGrade 9

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
5000/7 = 714 
714/7 = 102 
102/7 = 14 
14/7 = 2 

714+102+14+2 = 832 
 
Hence maximum power of 7 will be 832
 
 
Regards
Arun (askIITians forum expert)